# Theoretical study of orbital torque: Dependence on ferromagnet species and nonmagnetic layer thickness

### v1

@@ IV Conclusion
In this work, we examined current-induced torques in NM/FM bilayers using semi-realistic tight-binding models constructed from first-principles electronic structures. By directly evaluating the torque in heterostructures, our approach provides quantitative results without relying on a simplified picture based on the bulk properties of individual NM and FM layers. We systematically investigated the torque depending on material combination and NM layer thickness. Our results show that both Ti/FM and Cu/FM bilayers exhibit sizable damping-like torques, ascribed to the OAM generation in the bulk NMs. The dependence on the FM choice between Co and Ni, however, differs between Ti and Cu systems: Ni yields a larger torque in Ti/FM, whereas Co produces a stronger torque in Cu/FM. This contrasting behavior indicates that the FM-dependent OT efficiency is not governed solely by the SOC strength of the FM but is also strongly affected by orbital hybridization or interfacial characteristics across the NM/FM interface. Thickness-dependent analyses further reveal that the torque predominantly arises from the NM bulk, confirming the bulk nature of the OT. The sign of the OT does not necessarily coincide with that of the bulk orbital Hall conductivity, as also indicated by a recent experiment,Vijayan et al. (2025) calling for careful assessment of sign determination in OT measurements. The characteristic length ranges from the atomic scale to over 10 nm, depending on the disorder strength. These findings provide valuable guidance for engineering efficient OT devices based on light metals.

### v2

@@ IV Conclusion
In this work, we examined current-induced torques in NM/FM bilayers using realistic tight-binding models constructed from first-principles electronic structures. By directly evaluating the torque in heterostructures, our approach provides quantitative results without relying on a simplified picture based on the bulk properties of individual NM and FM layers. We systematically investigated the torque depending on material combination and NM layer thickness. Our results show that both Ti/FM and Cu/FM bilayers exhibit sizable damping-like torques, ascribed to the OAM generation in the bulk NMs. The dependence on the FM choice between Co and Ni, however, differs between Ti and Cu systems: Ni yields a larger torque in Ti/FM, whereas Co produces a stronger torque in Cu/FM. This contrasting behavior indicates that the FM-dependent OT efficiency is not governed solely by the SOC strength of the FM but is also strongly affected by orbital hybridization or interfacial characteristics across the NM/FM interface. Thickness-dependent analyses further reveal that the torque predominantly arises from the NM bulk, confirming the bulk nature of the OT. The sign of the OT does not necessarily coincide with that of the bulk orbital Hall conductivity, as also indicated by a recent experiment,Vijayan et al. (2025) calling for careful assessment of sign determination in OT measurements. The characteristic length ranges from the atomic scale to over 10 nm, depending on the disorder strength. These findings provide valuable guidance for engineering efficient OT devices based on light metals.